OpenStudy (anonymous):

f(x)=e^x+4lnx find f'(6).

OpenStudy (anonymous):

e^6 + 2/3

OpenStudy (anonymous):

show work please

OpenStudy (anonymous):

f'(x) = e^x + 4/x . d(lnx)/dx = 1/x so, f'(6) = e^6 + 2/3

OpenStudy (anonymous):

im getting 785.77 but thats wrong

OpenStudy (anonymous):

If I calculate the value of e^6 , it comes out to be approx 136.821. Then add 1.333 to it. Comes out to 138.154

OpenStudy (anonymous):

why am i not getting these numbers?

OpenStudy (anonymous):

you might be doing some silly mistake. Check again.

OpenStudy (anonymous):

when i type in e^6 i get 403.43

OpenStudy (anonymous):

in my calc

OpenStudy (anonymous):

nobody knows what we're doing wrong?

OpenStudy (turingtest):

\[e^6\approx403.4287\]according to my calculator

OpenStudy (anonymous):

well then I guess I am wrong...I'll check again! :P

OpenStudy (anonymous):

ditto

OpenStudy (turingtest):

oh I thought your decimal was in a different spot sidd

OpenStudy (turingtest):

but\[2.5^6=\frac{5^6}{2^6}=244.14...<e^6\]so mario's answer makes sense

OpenStudy (anonymous):

Yea....but if you see the expansion of e^x which is equal to 1 + x/1! + x^2/2! + x^3/3!.....so on, the value doesn't come out to be 403.287

OpenStudy (anonymous):

im so lost...

OpenStudy (kinggeorge):

Any way I look at it, I'm getting \(f'(6) = e^6 + {2 \over 3}\) Which is about 404.095

OpenStudy (anonymous):

thats incorrect george...

OpenStudy (kinggeorge):

Do you know what the answer is?

OpenStudy (anonymous):

Oh and its showing 403.428 on the calculator but why doesn't the same value come when I use the expansion of e^x?

OpenStudy (anonymous):

no. thats why im here. are we setting this up right?

OpenStudy (kinggeorge):

How do you know that's wrong then?

OpenStudy (anonymous):

i put it in my hw website and it tells me if i got it right or not

OpenStudy (kinggeorge):

Have you tried putting in the exact formula \(e^6 +{4 \over 6}\) ? Several of the websites I've used that have hw problems on them need it in the exact form, not decimal.

OpenStudy (anonymous):

it says, find f'(6), correct to two decimal places on there

OpenStudy (kinggeorge):

That's weird.

OpenStudy (anonymous):

Well mario I think the answer is 404.75. So just go with this.

OpenStudy (anonymous):

its not :C

OpenStudy (anonymous):

this is getting frustrating

OpenStudy (kinggeorge):

404.10? That's what I keep getting. If it's not that, I honestly have no idea what we're dong wrong.

OpenStudy (anonymous):

its not

OpenStudy (kinggeorge):

Ask your teacher. That's all I have left.

OpenStudy (anonymous):

Input: f (x) = e^x + 4 log (x)

OpenStudy (anonymous):

what is not going right? \[f'(x)=e^x +\frac{4}{x}\] \[f'(6)=e^6+\frac{2}{3}\]

OpenStudy (anonymous):

Juss said that :) ^

OpenStudy (anonymous):

404.09546015940178927505384721005494627256656402379586..

OpenStudy (anonymous):

Is the answer 397.43?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@mariomintchev do u have choices that can help us to help you..

OpenStudy (anonymous):

404.095.... rounds off to 404.10, right?

OpenStudy (anonymous):

its not multiple choice

OpenStudy (anonymous):

I don't know what the answer is then. Ask your teacher.

OpenStudy (anonymous):

thats what i mean. it says to round to 2 decimal places

OpenStudy (anonymous):

i gotta go for now but if u all come up with something, write it here.

OpenStudy (anonymous):

two decimal place accuracy gives 404.10

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=e%5E6%2B4%2F6

OpenStudy (anonymous):

e^6+4/6

OpenStudy (anonymous):

ok. im back. ive tried everything but i have been unsuccessful. any ideas? i can give you my hw website login info if you wanna try for yourself.

OpenStudy (turingtest):

it's not that hard of a problem the answer to the problem as you posted it is f'(6)=e^6+2/3 the possibilities are 1)your post has a typo 2)the website has a glitch

OpenStudy (turingtest):

\[f(x)=e^x+4\ln x\]\[f'(x)=e^x+\frac4x\]\[f'(6)=e^6+\frac23\]now if the problem is not this, then it makes sense otherwise that's all she wrote

OpenStudy (anonymous):

everything is raised

OpenStudy (anonymous):

not just the x

OpenStudy (anonymous):

e^(x+4lnx)

OpenStudy (anonymous):

Oh then the derivative is-- f'(x)=[ e^(x + 4lnx) ].[ 1 + 4/x] f'(6)= [ e^(6 + 2/3)].[ 1+ 2/3] No just calculate the value of that...